Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
amagazzu - 23:02 Sunday 11 February 2024 (63234) Print this report
Operator Report - Afternoon shift

The shift was dedicated to the planned SGD activity, Automation Squeezing Injection, carried out from remote by Zendri and Vardaro.
From 14:38 UTC to 15:15 UTC we manually realigned the FC cavity to lock it, Step READY_FOR_FIS_INJECTION on SQZ_MAIN reached at 15:15 UTC. At around 15:16 UTC we opened the SQB1 Shutter (-500000 steps) and start the injections.
The first sets of measurements were performed with the ITF in LOW_NOISE_2 (SR aligned state). Here the timetables of the measurements taken:

15:27 UTC - SQZ_INJECTED for 5 minutes;
15:33 UTC - SQZ_LOCKED for 10 minutes;
15:47 UTC - CC Phase scan (;

After this sets of measurements, we reached LOW_NOISE_3 at 16:15 UTC and we performed the rest of the tests (SR misaligned state):

16:34 UTC - SQZ_INJECTED for 5 minutes;
16:39 UTC - SQZ_LOCKED for 10 minutes;
16:50 UTC - CC Phase scan (;
17:37 UTC - Scan of the phase of the squeezing around the angle to maximize the range (;

The rest of the measurements consisted on swapping from SQZ_INJECTED to SQZ_LOCKED every 10 minutes. Here the timetables of the actions performed:


Activity still in progress, ITF left in LOW_NOISE_3 with AUTORELOCK_FAILSAFE Engaged.

Images attached to this report
Comments to this report:
zendri, vardaro - 22:50 Sunday 11 February 2024 (63235) Print this report
The CC phase scan (starting time 1391701795 duration 1268 sec.) in Low noise 2 configuration highlighted efficiency values in the range of typical values with the SR aligned (figure 1). However, it should be noted that the magnitude of the 4 MHz beat note is noisier than usual. The blue line in figure 2 represents the shot noise period (starting time 1391700836 duration 610) and the red line represents the phase scan data. The left plot shows that compared to the mean shot noise level at the squeezing angle we have a horizon increase slightly less than 2 Mpc

The phase scan in low noise 3 configuration is summarized in figures 3 and 4. Figure 3 shows that also in this case during the phase scan (starting time 1391705561 duration 1264 s ) the 4 MHz magnitude is noisy, while figure 4 shows a large variability of the horizon also in shot noise configuration (starting time 1391704750 duration 793 s). During the scan at around the 217 degree of the CC angle there seems to be a maximum of Hrec. For this reason we decide to do repeated scans around this maximum (from 195 to 235 degrees) which are summarized in figure 5. Despite the large fluctuations of Hrec these scans seem to highlight an increase in the horizon for a CC angle of approximately 3.85 radians. We then fix this angle and make on/off squeezing injections. Unfortunately, as evident in the figure 6, there does not appear to be an obvious effect of SQZ injection on Hrec. However, it is difficult to draw conclusions in this situation of great Hrec variability. The measurements will probably need to be repeated once sensitivity has stabilized and better checking the sqz alignment.
Images attached to this comment
vardaro, zendri, magazzu - 23:54 Sunday 11 February 2024 (63236) Print this report

At 20:59 Andrea put the automation in SQZ_INJECTED with the angle that maximized the range 2.7 rad.

After 10 min I relized that the B1 4MHz mag was only 1.4mV instead 1.8mV.

I decided to check the alignment of the system acting on SQB1_LC_X now at -3370 um and before at -3150um, then acting on the SQB1_RetroRefl H and V.

After some iterations I reached a B1_4MHz_mag of 1.92mV.

Then I decide to took 10 minutes of shot noise at 22:01 UTC.

At 22:11 I run again a CC phase scan (22:13 start of the scan)

At 22:48 I restarted to inject SQZ into the ITF. 

Note that:

in fig 2 you can see that during the phase scan the range was lower than before but the  b1 4 MHz mag had a maximum at 1.9mV

In fig 3 a comparison between hrec (no SQZ) during the on off and in the last part of the shift


We left the system in SQZ injected with the shutter open

Images attached to this comment
mwas - 8:02 Monday 12 February 2024 (63237) Print this report

Figure 1 the bump in the spectrum at ~35Hz started around 21:30 UTC. It is relatively broad, and is most likely scattered light from the squeezing, involving ground motion which makes a 35Hz mechanical resonance of the scattering surface modulated by the ground motion look like a broad peak instead of a narrow resonance.

Images attached to this comment
direnzo - 20:44 Monday 12 February 2024 (63243) Print this report

To investigate the origin of the excess noise at low frequency highlighted in Michal's entry, I estimated the BLRMS of Hrec in the band 30 and 40 Hz and correlated it with all trend *_rms channels (~9k). The BLRMS is estimated with the modified algorithm described in VIR-0484A-21. This estimate automatically ignores spectral lines, which would dominate the RMS value in the band, and glitches, which would influence the Pearson correlation.

Figure 1: The most correlated channels are those in the V1:SQB1_LC_LVDT_{BL,FL,BR,FR}_H_out_raw_DS_rms family.

I then repeated the analysis to include all the ENV_* and SQZ_* channels (~2k). In the attached figures I report those that look most interesting.

Figure 2: V1:SQZ_AA_SLOW_Cam{0,1}Y_ERR

Figure 3: V1:ENV_DT_CT_FINGER_ACC_Y_min

Figure 4: V1:SQZ_MIRROR2_X_cmd_1kHz_*

Images attached to this comment
vardaro - 11:29 Wednesday 14 February 2024 (63271) Print this report

I add some details. I used as figure of merit to improve the alignment between SQZ and ITF the B1 4 MHz mag, i.e. the beat note between the CC field and the ITF beam.

As you can see I increased it between 21 and 21:30 UTC acting both

  • on the SQB1 Retroreflector X and Y 
  • on SQB1 LC X

You can see from fig 1 that the 35 Hz bump started after the movement of SQB1 LC X. I reverted its position on Monday morning

Please note that moving the retroreflector can also move the position of the beam on SQB1 AA mirror increasing the scattered light (to be checked)

In addition on Monday I checked the mag of the 4 MHz on the EQB1 HD Detector and I found a lot of losses. I had to move back the retroreflector to decrease this amount of losses. I don't understand why the mag into the ITF increased.


Images attached to this comment
mwas - 12:15 Wednesday 14 February 2024 (63272) Print this report

Figure 1. There are angular dither lines in the squeezing beam path, one of them is at 35Hz. So the bump at 35Hz could be related to that particular angular dithering.

Images attached to this comment
Search Help